OpenStudy (anonymous):

75*803=? pzz help show answer

OpenStudy (mathsolver):

60225

OpenStudy (anonymous):

plzz help

OpenStudy (mathsolver):

what do you need?

hero (hero):

Hmmmm

hero (hero):

She might need a more in-depth explanation. You should have asked her what grade she's in first.

OpenStudy (anonymous):

sixth

OpenStudy (anonymous):

If I had to do it without a calculator: \[75 = 25\cdot 3\]\[803 = 800 + 3 = 400\cdot 2 + 3 = 100\cdot 4\cdot 2 + 3\]\[\therefore75\cdot 803 \]\[= (25\cdot3)(100\cdot 4\cdot 2 + 3)\]\[=25\cdot 4\cdot 100\cdot 2\cdot 3 + 25\cdot3\cdot 3\]\[=100\cdot 100\cdot 6 + 25\cdot 9\]\[=60000 + 225 \]\[= 60225\]

OpenStudy (anonymous):

u welcom 4 the medels

OpenStudy (anonymous):

u welcom 4 the medels

hero (hero):

Thanks...I didn't expect it at all

OpenStudy (anonymous):

But you can also do it with 'long' multiplication: 803 x 75 ====== 4015 + 56210 ====== 60225

hero (hero):

That looks very nice. I like the top down method. So organized and simple.

OpenStudy (anonymous):

kk thank u

hero (hero):

Oh and Polpak, I've been taking your advice....

OpenStudy (anonymous):

oh?

OpenStudy (anonymous):

Which advice is that?